a manufacture has 15 years, 8.5 percent mortgage for 100,000 birr on a building. payments are made monthly.

a. find the monthly payment

b. how much interest will be paid?


1
Expert's answer
2022-03-04T09:23:00-0500

a.

The formula for monthly payment is:

Fixed monthly payment ="\frac{P(r/n)(1+r/n)^{nt}}{(1+r/n)^{nt}-1}"

Where Principal P = 100,000 birr

For 15-year mortgage at 8.75%

n=12months per year, t=15 years, rate r=0.0875 yearly

Therefore,

The monthly payment = "\frac{P(r/n)(1+r/n)^{nt}}{(1+r/n)^{nt}-1}"

= "\frac{100,000(\frac{0.0875}{12})(1+\frac{0.0875}{12})^{180}}{(1+\frac{0.0875}{12})^{180}-1}"

= 999.45 birr

Answer: Monthly payment = 999.45 birr


b.

For the 15-year fixed mortgage at 8.75%, the monthly payment = 999.45 birr for (12x15) = 180 periods

Therefore, the total amount to be paid = (999.45 x 180)= 179,901 birr

The interest = total amount – principal = (179,901- 100,000) = 79,901 birr


Answer: The interest to be paid = 79,901 birr